### v2

# Hyperspectral Image Data Reduction for Endmember Extraction

Abstract
Endmember extraction from hyperspectral images aims to identify the spectral signatures of materials present in a scene. Recent studies have shown that self-dictionary methods can achieve high extraction accuracy; however, their high computational cost limits their applicability to large-scale hyperspectral images. Although several approaches have been proposed to mitigate this issue, it remains a major challenge. Motivated by this situation, this paper pursues a data reduction approach. Assuming that the hyperspectral image follows the linear mixing model with the pure-pixel assumption, we develop a data reduction technique that removes pixels that do not contain endmembers. We analyze the theoretical properties of this reduction step and show that it preserves pixels that lie close to the endmembers. Building on this result, we propose a data-reduced self-dictionary method that integrates the data reduction with a self-dictionary method based on a linear programming formulation. Numerical experiments demonstrate that the proposed method can substantially reduce the computational time of the original self-dictionary method without sacrificing endmember extraction accuracy.
Keywords: hyperspectral image, endmember extraction, self-dictionary method, data reduction, nonnegative matrix factorization
1 Introduction
Endmember extraction from a hyperspectral image (HSI) is a fundamental task in utilizing hyperspectral imaging technology for Earth’s surface observation, with applications ranging from vegetation monitoring to mineral and resource exploration. Numerous endmember extraction methods have been proposed to date; see [5, 22, 19] for surveys of recent advances. Among these, self-dictionary methods have received growing attention because of their ability to achieve high extraction accuracy on HSIs. However, a major drawback of self-dictionary methods is their high computational cost, which hinders their applicability to large-scale HSIs.
Self-dictionary methods formulate the endmember extraction task as a sparse optimization problem, treating the HSI data itself as a dictionary, and then solve a convex relaxation of this problem. Endmembers are identified using the optimal solution of the relaxed problem. Although various self-dictionary methods have been developed in the literature [6, 10, 11, 15, 27], this paper mainly focuses on the method proposed by Bittorf et al. [6], which employs a linear programming (LP) formulation as the convex relaxation. For brevity, we refer to this LP-based self-dictionary method simply as the “LP method” throughout the paper. Solving this LP relaxation is the primary computational bottleneck in applying the LP method to large-scale HSIs, because the size of the LP problem grows quadratically with the number of pixels in the image.
There are two main approaches to addressing this computational issue: (i) developing efficient algorithms for solving the convex relaxation problem, and (ii) developing preprocessing techniques that reduce the size of the HSI data before applying self-dictionary methods. There is a line of research exploring the first approach. For example, an efficient algorithm based on the column generation framework was proposed in [25] for the LP method. However, even with this improvement, processing the Urban HSI dataset, which is a widely used benchmark in the remote sensing community, still takes 8 hours. In light of this situation, this paper pursues the second approach.
Signal Model for Endmember Extraction from HSIs
To describe our research question and contributions, we begin by introducing the signal model for endmember extraction from HSIs. In this paper, we adopt the linear mixing model under the pure-pixel assumption. Consider an HSI consisting of $n$ pixels, acquired by a hyperspectral sensor with $d$ spectral bands. We represent the HSI as a matrix $A\in\mathbb{R}^{d\times n}$, where the $(i,j)$th entry denotes the measurement of the $i$th spectral band at the $j$th pixel. Thus, the $j$th column of $A$ corresponds to the observed spectral signature at the $j$th pixel. We refer to such a matrix $A$ as an HSI matrix. Given an HSI matrix $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$, a linear mixing model, abbreviated as LMM, assumes that the observed spectral signatures $\bm{a}_{1},\ldots,\bm{a}_{n}$ are generated as
$\bm{a}_{i}=\sum_{j=1}^{r}h_{ji}\bm{w}_{j}+\bm{v}_{i}\quad\text{for}\ i=1,\ldots,n,$
where each $\bm{w}_{j}$ satisfies $\bm{w}_{j}\geq\bm{0}$ and each $h_{ji}$ satisfies $\sum_{j=1}^{r}h_{ji}=1$ and $h_{ji}\geq 0$. We refer to $\bm{w}_{j}$ as an endmember signature, $h_{ji}$ as the abundance fraction of the $j$th endmember in the $i$th pixel, and $\bm{v}_{i}$ as the noise associated with the $i$th pixel. An endmember signature represents the spectral signature of a distinct material present in the observed scene, while an endmember refers to the material itself.
To explain the pure-pixel assumption, we express the model above in matrix form as
$A=WH+V,$ (1)
where $W=[\bm{w}_{1},\ldots,\bm{w}_{r}]\in\mathbb{R}^{d\times r},H\in\mathbb{R}^{r\times n}$ with $(j,i)$th element $h_{ji}$, and $V=[\bm{v}_{1},\ldots,\bm{v}_{n}]\in\mathbb{R}^{d\times n}$. Note that both $W$ and $H$ are nonnegative matrices, and, in particular, the sum of the elements in each column of $H$ equals one. We refer to $W$ as the endmember matrix of $A$. We say that the $j$th endmember has a pure pixel if there exists a column of $H$ equal to the $j$th unit vector. The pure-pixel assumption asserts that each endmember has at least one such pure pixel. Equivalently, $H$ can be written in the form
$H=[I,\bar{H}]\Pi\in\mathbb{R}^{r\times n},$ (2)
where $I$ is the $r\times r$ identity matrix, $\Pi$ is an $n\times n$ permutation matrix, and $\bar{H}$ is an $r\times(n-r)$ submatrix. When the pure-pixel assumption holds, the LMM implies that, in the absence of noise (i.e., $V=0$), the endmember signatures $\bm{w}_{1},\ldots,\bm{w}_{r}$ appear exactly as columns of $A$.
Motivation and Research Question
We examine an HSI matrix $A$ from a geometric perspective, assuming that it follows the LMM with the pure-pixel assumption, as given in (1) and (2). Let $\operatorname{cone}(A)$ denote the conical hull of the columns of $A$. To simplify the discussion, we assume $V=0$. In this setting, the endmember signatures $\bm{w}_{1},\ldots,\bm{w}_{r}$ correspond to the extreme rays of $\operatorname{cone}(A)$. Thus, by removing the columns of $A$ that lie in its interior, the extreme rays remain, which correspond to the endmembers. When $V\neq 0$, the resulting columns may instead be close to the endmembers. Determining whether a column of $A$ lies in the interior of $\operatorname{cone}(A)$ can be formulated as a convex programming problem, and the associated computational cost is low. Accordingly, if this data reduction step succeeds in discarding only irrelevant columns while retaining those close to the endmembers, it can substantially reduce the computational time of self-dictionary methods without compromising their extraction performance.
Figure 1 illustrates this observation using the Samson HSI dataset, which is described in detail in Section 5.1. Hereafter, let $A$ denote the HSI matrix of the Samson dataset. We computed the intersection points of the columns of $A$ with a hyperplane and then projected them onto a two-dimensional space. The figure shows the projected points of all columns of $A$ (left) and those obtained after removing the columns lying in the interior of $\operatorname{cone}(A)$ (right). It can be seen that a large number of columns of $A$ are removed without discarding columns close to the endmembers. Specifically, $A$ has 9025 columns, but after removing the columns contained in the interior of $\operatorname{cone}(A)$, only 20 columns remain. The details of how this figure was generated are provided in Section 5.3.
At this point, a natural question arises: After removing the columns that lie in the interior of the conical hull of the HSI matrix, how close are the remaining ones to the endmembers? To the best of our knowledge, this research question has not been investigated from a theoretical perspective in the existing literature. In the works of Esser et al. [11] and Gillis and Luce [15], data reduction techniques based on clustering were employed within self-dictionary methods. However, the theoretical properties of these reduction steps were not analyzed in their studies. In this paper, we address this question through a theoretical analysis. Building on the resulting insights, we then develop a data reduction algorithm, which we integrate into a self-dictionary method.
Figure 1: Example of the Samson HSI dataset. The black dots represent the columns of the HSI matrix $A$, while the red x-markers represent the columns of the endmember matrix $W$. The left panel shows all columns of $A$, and the right panel shows the columns of AA after removing those lying in the interior of $\operatorname{cone}(A)$.
Contributions
The main contributions of this paper are threefold. First, we address the research question posed above. Our theoretical analysis is conducted under the assumption that the HSI matrix follows the LMM with the pure-pixel assumption. We summarize our result as Theorem 1, showing that even after removing the columns located inside the conical hull of the HSI matrix, there still exist columns that lie close to the endmembers, provided that the noise level is below a certain threshold.
Second, we develop a data reduction algorithm based on the theoretical result, and subsequently propose a data-reduced self-dictionary method, abbreviated as REDIC, for endmember extraction from HSIs by integrating data reduction with the variant of the LP method proposed in [25]. As observed in the case of the Samson dataset, applying data reduction to HSI matrices can remove a large portion of their columns. However, directly applying the LP method to the reduced data does not necessarily yield high extraction accuracy. To address this issue, we incorporate a data augmentation process together with an averaging technique to improve the extraction accuracy of REDIC. For the reduced matrix obtained after data reduction, we augment it by randomly selecting some of the removed columns and adding them back. This augmentation step helps improve the accuracy of the extracted endmembers. However, the randomness involved in the augmentation process introduces variability into the extraction results. To mitigate this variability, we repeat the augmentation and endmember extraction using the LP method multiple times. Finally, we average the outputs produced by the LP method to obtain the final estimates of the endmembers.
Third, we present numerical experiments demonstrating the performance of REDIC. The experiments were conducted using both synthetic and real HSI datasets. The results indicate that the data reduction algorithm within REDIC can remove a large portion of the columns in HSI matrices; for example, for the Urban dataset, which originally contains about one hundred thousand columns, only 483 columns remain after applying this data reduction. In addition, among the remaining columns, there exists at least one column close to each of the endmembers. We compared the endmember extraction accuracy and computation time of REDIC with those of the LP method in [25]. The results show that REDIC can achieve extraction accuracy comparable to that of the LP method while being substantially faster; for the Urban dataset, REDIC requires only 16 minutes, whereas the LP method takes 8 hours.
The remainder of this paper is organized as follows. Section 2 formulates the endmember extraction problem and reviews self-dictionary methods. Section 3 presents our theoretical result on data reduction and describes the data reduction algorithm. Section 4 describes the REDIC method in detail. Section 5 reports numerical experiments. Section 6 provides the proof of the theoretical result.
Notation
We write $\mathbb{R}^{d\times n}_{+}$ for the set of nonnegative matrices of size $d\times n$. We use $\bm{0}$ to denote the vector of all zeros, $\bm{1}$ the vector of all ones, $\bm{e}_{i}$ the $i$th unit vector, $O$ the matrix of all zeros, $I$ the identity matrix, and $\Pi$ a permutation matrix. For a positive integer $n$, we write $n]=\{1,\ldots,n\}$. For a set $\mathcal{S}\subset[n]$ and an element $s\in\mathcal{S}$, we use $\mathcal{S}-s$ to denote the set obtained by removing $s$ from $\mathcal{S}$, i.e., $\mathcal{S}\setminus\{s\}$.
Let $\bm{a}\in\mathbb{R}^{d}$ and $A\in\mathbb{R}^{d\times n}$. The notation $\bm{a}(i)$ denotes the $i$th entry of $\bm{a}$, and $A(i,j)$ denotes the $(i,j)$th entry of $A$. We denote by $A^{\top}$ the transpose of $A$ and by $\operatorname{rank}(A)$ the rank of $A$. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]$. For $\mathcal{K}\subset[n]$, we use $A(\mathcal{K})$ to denote the submatrix of $A$ obtained by retaining only the columns whose indices belong to $\mathcal{K}$. For example, if $\mathcal{K}=\{1,3\}$, then $A(\mathcal{K})=[\bm{a}_{1},\bm{a}_{3}]$. The conical hull of the columns of $A$ is denoted by $\operatorname{cone}(A)$, i.e., $\operatorname{cone}(A)=\{A\bm{x}\mid\bm{x}\geq\bm{0}\}$. We use $\|\cdot\|$ to denote a norm on a vector or matrix. In particular, $\|\cdot\|_{p}$ denotes the $L_{p}$ norm, and $\|\cdot\|_{F}$ the Frobenius norm.
2 Self-Dictionary Approaches to the Endmember Extraction Problem
2.1 Problem Formulation
We define the endmember extraction problem as described in Problem 1, where we assume that the HSI matrix follows the LMM under the pure-pixel assumption and that the number of endmembers is known in advance.
Problem 1.
Suppose that an HSI matrix $A\in\mathbb{R}^{d\times n}$ is generated by the LMM $A=WH+V$, as given in (1), where the pure-pixel assumption holds; that is, $H$ has the form (2). Given the HSI matrix $A$ and the number of endmembers $r$, the objective is to estimate the $r$ columns of the endmember matrix $W$.
In this paper, we use the mean-removed spectral angle (MRSA) to evaluate the similarity between the true endmember signatures and their estimates. For $\bm{c}\in\mathbb{R}^{d}$, we define $\mathrm{ave}(\bm{c})=(\bm{1}^{\top}\bm{c}/d)\cdot\bm{c}$. The MRSA value between the spectral signature vectors $\bm{a},\bm{b}\in\mathbb{R}^{d}$ is defined as
$\mathrm{MRSA}(\bm{a},\bm{b})=\frac{1}{\pi}\arccos\frac{(\bm{a}-\mathrm{ave}(\bm{a}))^{\top}(\bm{b}-\mathrm{ave}(\bm{b}))}{\|\bm{a}-\mathrm{ave}(\bm{a})\|_{2}\,\|\bm{b}-\mathrm{ave}(\bm{b})\|_{2}}\in[0,1].$
A smaller MRSA value for $\bm{a}$ and $\bm{b}$ indicates that $\bm{a}$ is more similar to $\bm{b}$.
Problem 1 is essentially equivalent to computing nonnegative matrix factorizations (NMFs) under the separability assumption. Considerable research has been devoted to the separable NMF problem; see [12, 19] for surveys of recent advances. Arora et al. [2] showed that $W$ can be recovered from $A$ in polynomial time when $V=O$. Even when $V\neq O$, it is still possible to extract a submatrix of $A$ that is close to $W$, provided that the noise $V$ is below a certain threshold. Separable NMFs have applications not only to endmember extraction from HSIs, but also to topic modeling in document corpora [4, 3]. Other applications include community detection [20], spectral clustering [23], and crowdsourcing [21].
2.2 Review of Self-Dictionary Methods
Numerous endmember extraction methods have been proposed to date. These methods can be broadly classified into two categories: greedy methods and convex programming-based methods. Representative examples of greedy methods include PPI [7], N-FINDR [28], SPA [1, 16], and VCA [26], all of which are widely used for endmember extraction. Greedy methods have low computational cost and can therefore handle large-scale datasets, although the accuracy of the extracted endmembers may not always be sufficient. In contrast, convex programming-based methods generally achieve higher extraction accuracy, but their high computational cost remains a drawback.
We consider self-dictionary approaches within the category of convex programming-based methods. Let $A\in\mathbb{R}^{d\times n}$ denote an HSI matrix, and let $r$ denote the number of endmembers. Regarding the matrix $A$ as a dictionary and its columns as atoms, we interpret Problem 1 as the task of identifying $r$ atoms from this dictionary such that the entire dictionary can be well approximated by a convex combination of the selected atoms. Based on this interpretation, we formulate the problem as the following sparse optimization problem:
$\min_{X\in\mathbb{R}^{n\times n}}$ $\|A-AX\|$ subject to $\|X\|_{\mathrm{row},0}=r,\ \bm{1}^{\top}X=\bm{1}^{\top},\ X\geq O,$ (3)
where $\|X\|_{\mathrm{row},0}$ denotes the number of nonzero rows of $X$.
Solving problem (3) exactly is difficult because of the combinatorial constraint $\|X\|_{\mathrm{row},0}=r$. Self-dictionary methods therefore typically rely on a convex relaxation of this problem. As mentioned in Section 1, this paper focuses on the LP method proposed by Bittorf et al. [6], which employs an LP formulation as the convex relaxation. This method was originally developed in the context of identifying hot topics in document corpora and was therefore named Hottopixx. Since its original proposal, the LP method has been extended and refined by several researchers [17, 14, 24]. In what follows, we review the version of the LP method proposed in [24]. This method is based on the following optimization model:
$\mathsf{H}:$ $\min_{X\in\mathbb{R}^{n\times n}}$ $\|A-AX\|_{1}$
subject to $\sum_{i=1}^{n}X(i,i)=r,$
$0\leq X(i,j)\leq X(i,i)\leq 1,$ $i,j=1,\ldots,n.$
Model $\mathsf{H}$ relaxes the combinatorial constraint $\|X\|_{\mathrm{row},0}=r$ in problem (3) and can therefore be reformulated as an LP. Feasible solutions $X$ may have more than $r$ nonzero rows. However, the number of such rows is expected to be small, since the constraints restrict the number of nonzero diagonal elements in $X$ and, moreover, if the $i$th diagonal element $X(i,i)$ is zero, then the $i$th row of $X$ must also be zero.
The LP method consists of two main steps: it first computes the optimal solution $X_{\mathrm{opt}}$ of $\mathsf{H}$ and then performs a postprocessing step based on $X_{\mathrm{opt}}$. In the postprocessing step, $r$ clusters of the column indices of the input matrix $A$ are formed according to the diagonal elements of $X_{\mathrm{opt}}$. One representative element is then chosen from each cluster and returned as the output. Theorem 3.2 of [24] guarantees that the corresponding columns of $A$ indexed by these elements are close to the $r$ endmember signatures, provided that the noise level is below a certain threshold.
From a theoretical perspective, the LP method is attractive, as the theoretical result suggests that the accuracy of the extracted endmembers can be high. However, its major drawback lies in the high computational cost. Specifically, the method requires solving the model $\mathsf{H}$, which reduces to an LP with $O(n^{2})$ variables and $O(n^{2})$ constraints; see [25] for details. Consequently, the method becomes impractical for HSIs with a large number of pixels $n$. To address this computational issue, a row and column expansion (RCE) algorithm was proposed in [25] to solve model $\mathsf{H}$ efficiently. The RCE algorithm is based on a column generation technique, a well-known approach for solving large-scale LPs. Nevertheless, even with RCE, the improvement in computation time is insufficient. For example, as shown in Section 5.3, the Urban dataset requires 8 hours of computation, whereas greedy methods such as SPA and VCA finish in only a few seconds. This observation motivates us to develop a data reduction technique for hyperspectral images.
In addition to the LP method, various self-dictionary approaches have been proposed for endmember extraction from HSIs [10, 11, 15, 27]. FGNSR, proposed by Gillis and Luce [15], and MERIT, proposed by Nguyen et al. [27], are most closely related to the LP method. These methods are based on the following optimization model:
$\min_{X\in\mathbb{R}^{n\times n}}$ $\frac{1}{2}\|A-AX\|_{F}^{2}+\lambda\cdot\Phi(X)$ subject to $X\in\mathcal{F}$
where $\lambda$ is a penalty parameter and $\Phi$ is a regularization term designed to promote row sparsity in $X$. FGNSR and MERIT identify endmembers by leveraging the optimal solution of this model. As with the LP method, solving this model incurs a high computational cost and becomes impractical for HSIs with a large number of pixels. The work by Esser et al. [11] can be viewed as a self-dictionary approach to the endmember extraction problem, whereas the work by Elhamifar et al. [10] addresses a related problem within a similar modeling framework.
3 Theoretical Result on Data Reduction and Algorithm Development
3.1 Main Theorem
We now present the main theoretical result (Theorem 1) that addresses the research question posed in Section 1. To this end, we introduce the necessary terminology. Let $M\in\mathbb{R}^{d\times n}_{+}$ be factored as
$M=WH\in\mathbb{R}^{d\times n}_{+},\quad W\in\mathbb{R}^{d\times r}_{+},\quad H=[I,\bar{H}]\Pi\in\mathbb{R}^{r\times n}_{+}.$ (4)
We say that such a matrix $M$ is $r$-separable. If an $r$-separable matrix $M$ is perturbed by noise so that $A=M+V$ for $V\in\mathbb{R}^{d\times n}$, we say that $A$ is nearly $r$-separable. Accordingly, if an HSI matrix follows the LMM with the pure-pixel assumption, it is nearly $r$-separable. Our analysis makes the following assumption regarding the nearly $r$-separable matrix $A=M+V$:
Assumption 1.
Let $A=M+V\in\mathbb{R}^{d\times n}$ be a nearly $r$-separable matrix, where $M\in\mathbb{R}^{d\times n}_{+}$ is $r$-separable of the form $M=WH$, as shown in (4), and $V\in\mathbb{R}^{d\times n}$ is noise. We assume that the following conditions hold:
  1. (a)
    Each column of $W$ and $H$ has unit $L_{1}$ norm.
  2. (b)
    $\|V\|_{1}\leq\epsilon$ for some $\epsilon\geq 0$.
A similar assumption has been used in the analyses of the noise robustness of LP methods for separable NMFs [17, 24]. While [24] assumes $0\leq\epsilon<1$, this condition is automatically satisfied in the setting of Theorem 1 due to the remaining assumptions. Part (a) can be imposed without loss of generality for $r$-separable matrices. Our analysis employs the conditioning parameter $\rho(W)$ for $W\in\mathbb{R}^{d\times r}_{+}$, defined by
$\rho(W)=\min_{\|\bm{x}\|_{1}=1}\|W\bm{x}\|_{1}.$
This parameter serves as a measure of the linear independence of the columns of $W$. In particular, $\rho(W)>0$ if and only if the $r$ columns of $W$ are linearly independent. Moreover, if part (a) of Assumption 1 holds, then $\rho(W)\leq 1$. These properties are formally established in Lemmas 1 and 2. Since our analysis relies on the linear independence of the columns of $W$, we make use of the parameter $\rho(W)$. For $A\in\mathbb{R}^{d\times n}$, define the family $\Gamma(A)$ of $\mathcal{K}\subset[n]$ as
$\Gamma(A)=\{\mathcal{K}\subset[n]\mid\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}))\}.$
Our theoretical result is summarized in the following theorem.
Theorem 1.
Let $A\in\mathbb{R}^{d\times n}$ be a nearly $r$-separable matrix of the form $A=M+V$, where $M=WH$ is $r$-separable and $V$ is noise. Assume that Assumption 1 holds for $A$. Let $\mathcal{K}\in\Gamma(A)$. If $\epsilon<\rho(W)/9$, then there exist $k_{1},\ldots,k_{r}\in\mathcal{K}$ such that
  • $k_{1},\ldots,k_{r}$ are all distinct, and
  • $\|\bm{w}_{j}-\bm{a}_{k_{j}}\|_{1}<(9/\rho(W)+1)\epsilon$ for each $j\in[r]$.
The proof is given in Section 6.
3.2 Data Reduction Algorithms
We develop algorithms to find a set $\mathcal{K}\in\Gamma(A)$ with a small number of elements. Algorithm 1, referred to as DR, describes a procedure for identifying a small set $\mathcal{K}\in\Gamma(A)$.
Algorithm 1 DR: Data reduction
Input: $A\in\mathbb{R}^{d\times n}$
Output: $\mathcal{K}\subset[n]$
  1. 1.
    Initialize $\mathcal{K}_{1}\leftarrow[n]$ and $\ell\leftarrow 1$.
  2. 2.
    For $i=1,\ldots,n$, do: if $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$, then set $\mathcal{K}_{\ell+1}\leftarrow\mathcal{K}_{\ell}-i$ and increment $\ell\leftarrow\ell+1$.
  3. 3.
    Set $\mathcal{K}\leftarrow\mathcal{K}_{\ell}$ and return $\mathcal{K}$.
DR generates a finite number of sets $\mathcal{K}_{1},\ldots,\mathcal{K}_{t}$ that satisfy $\operatorname{cone}(A(\mathcal{K}_{1}))=\cdots=\operatorname{cone}(A(\mathcal{K}_{t}))$. In particular, since $\mathcal{K}_{1}=[n]$ and $\mathcal{K}_{t}=\mathcal{K}$, the final output $\mathcal{K}$ belongs to $\Gamma(A)$. Moreover, removing any element from $\mathcal{K}$ yields a set that no longer belongs to $\Gamma(A)$. Therefore, the output $\mathcal{K}$ is a set with the smallest number of elements in $\Gamma(A)$. A formal statement of this result is provided in the Appendix.
To perform the feasibility test of whether $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$, we formulate it as a nonnegative least-squares problem:
$\min_{\bm{x}\geq\bm{0}}\quad\|A(\mathcal{K}_{\ell}-i)\bm{x}-\bm{a}_{i}\|_{2}^{2}.$ (5)
Efficient algorithms for solving this problem are available, such as the active set method and the projected gradient method. The optimal value is zero if and only if $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$. We adopt this formulation for the feasibility test in Step 2. Let $\bm{x}_{\mathrm{sol}}$ denote the numerically obtained solution to problem (5). If $\bm{x}_{\mathrm{sol}}$ satisfies
$\|A(\mathcal{K}_{\ell}-i)\bm{x}_{\mathrm{sol}}-\bm{a}_{i}\|_{2}<\epsilon_{\mathrm{feas}},$
where $\epsilon_{\mathrm{feas}}$ is a small positive tolerance, we regard $\bm{a}_{i}$ as belonging to $\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$ and remove $\bm{a}_{i}$ from $A$.
To enhance the computational efficiency of DR, we employ a data-splitting technique. To simplify the description of this procedure, we introduce the following notation. For a positive integer $m$, let $\mathcal{I}$ be a set of $m$ positive integers and $\mathcal{J}$ be a subset of $[m]$. We use the notation $\mathcal{I}(\mathcal{J})$ to denote the set $\{i_{j}\mid j\in\mathcal{J}\}$, where $i_{1},\ldots,i_{m}$ are the elements of $\mathcal{I}$ arranged in ascending order. The DR algorithm with data splitting proceeds as follows. First, we cluster the $n$ columns of $A$ into $p$ groups using the $k$-means method. Let $\mathcal{I}_{1},\ldots,\mathcal{I}_{p}$ denote the resulting $p$ groups of the column index set $[n]$. These groups satisfy $\mathcal{I}_{1}\cup\cdots\cup\mathcal{I}_{p}=[n]$ and $\mathcal{I}_{i}\cap\mathcal{I}_{j}=\emptyset$ for all distinct $i,j\in[p]$. The collection $\{\mathcal{I}_{1},\ldots,\mathcal{I}_{p}\}$ is referred to as a $p$-way partition of $[n]$. Next, we apply DR to each submatrix $A(\mathcal{I}_{u})$ for $u\in[p]$ and obtain the corresponding outputs, denoted by $\mathcal{J}_{1},\ldots,\mathcal{J}_{p}$. For each $u\in[p]$, set $\mathcal{K}_{u}=\mathcal{I}_{u}(\mathcal{J}_{u})$. We then construct $\mathcal{I}=\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}$ and apply DR once more to the submatrix $A(\mathcal{I})$. Let $\mathcal{J}$ be the output, and set $\mathcal{K}=\mathcal{I}(\mathcal{J})$. We return $\mathcal{K}$ as the final output. The overall procedure is summarized in Algorithm 2, which we refer to as DRS. Similar to DR, the output $\mathcal{K}$ of DRS is a set with the smallest number of elements in $\Gamma(A)$. A formal statement of this result is provided in the Appendix.
Algorithm 2 DRS: Data reduction via splitting
Input: $A\in\mathbb{R}^{d\times n}$; parameter $p$
Output: $\mathcal{K}\subset[n]$
  1. 1.
    Apply the $k$-means method to the columns of $A$ to construct the $p$-way partition $\{\mathcal{I}_{1},\ldots,\mathcal{I}_{p}\}$ of the column index set $[n]$ of $A$.
  2. 2.
    For $u=1,\ldots,p$, apply DR to $A(\mathcal{I}_{u})$ to obtain the output $\mathcal{J}_{u}$. Set $\mathcal{K}_{u}=\mathcal{I}_{u}(\mathcal{J}_{u})$.
  3. 3.
    Let $\mathcal{I}=\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}$. Apply DR to $A(\mathcal{I})$ to obtain the output $\mathcal{J}$. Set $\mathcal{K}=\mathcal{I}(\mathcal{J})$ and return $\mathcal{K}$.
We conclude this section with a discussion of related work. Finding a set $\mathcal{K}\in\Gamma(A)$ with the smallest number of elements is referred to as the conical hull problem in the paper by Dulá et al. [9]. They proposed an efficient algorithm for solving this problem and reported numerical results. Starting with $\mathcal{K}=\emptyset$, their algorithm constructs $\mathcal{K}$ by sequentially adding elements. This stands in contrast to our DR algorithm, which begins with $\mathcal{K}=[n]$, where $n$ is the number of columns of $A$, and removes elements one by one. At each iteration, their algorithm performs a feasibility test using an LP formulation, whereas our DR algorithm employs a nonnegative least-squares formulation. As mentioned in [13], their algorithm can be regarded as an instance of the algorithmic framework studied in [8].
4 The REDIC Method for Endmember Extraction
We develop REDIC, a data-reduced self-dictionary method for endmember extraction from HSIs by combining DRS with the LP method of [24]. Here, we describe the details of REDIC. To implement the LP method, REDIC employs the approach proposed in [25]. This approach is referred to as EEHT (efficient and effective Hottopixx), where Hottopixx is the original name of the LP method.
First, EEHT applies a dimensionality reduction technique to the input HSI matrix. Given $A\in\mathbb{R}^{d\times n}$ and a positive integer $r$, we compute the top-$r$ truncated singular value decomposition (SVD) of $A$ as
$A_{r}=U_{r}\Sigma_{r}V_{r}^{\top}.$
Here, $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix whose entries are the top $r$ singular values of $A$, and $U_{r}\in\mathbb{R}^{d\times r}$ (resp. $V_{r}\in\mathbb{R}^{n\times r}$) are the left (resp. right) singular vectors associated with these singular values. We then construct a reduced matrix $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$. It should be noted that if $A$ follows the LMM, i.e., $A=WH+V$, then $A^{\prime}$ also follows the LMM, i.e., $A^{\prime}=W^{\prime}H+V^{\prime}$, but with $r$ rows instead of $d$ rows. Second, EEHT constructs the model $\mathsf{H}$ using $A^{\prime}$ instead of $A$ and solves it with the RCE algorithm proposed in [25]. Finally, EEHT performs a postprocessing step: it partitions the column index set of $A$ into $r$ clusters based on the optimal solution of $\mathsf{H}$ and chooses elements from these clusters according to a prescribed criterion. For a detailed description of EEHT, we refer readers to Algorithm 4 in [25]. The experimental results demonstrate that RCE significantly reduces the computation time required to solve $\mathsf{H}$. Moreover, the accuracy of the extracted endmembers can be improved when the postprocessing step chooses representative elements from clusters based on their centroids, a procedure referred to as method-C. EEHT employing method-C for postprocessing is denoted as EEHT-C in [25].
We design the framework of REDIC as follows. First, it applies SVD-based dimensionality reduction and DRS to the input matrix as a preprocessing step and then constructs a reduced matrix. Second, it builds the model $\mathsf{H}$ using the reduced matrix and solves it with RCE. Finally, it performs the postprocessing using method-C. As shown in Section 5, this preprocessing step allows for the elimination of a large number of columns from HSI matrices. However, the performance of endmember extraction achieved by this framework is not always satisfactory. Therefore, to improve its performance, we incorporate a data augmentation technique and an averaging technique into the framework.
Let $A^{\prime}$ be the reduced matrix obtained by applying SVD-based dimensionality reduction to the input matrix $A$. Let $\mathcal{K}$ be the index set obtained after applying DRS to $A^{\prime}$. The data augmentation technique randomly selects a predetermined number of elements from the set $n]\setminus\mathcal{K}$ and adds them to $\mathcal{K}$. After this augmentation, endmember extraction is performed on the submatrix of $A^{\prime}$ whose columns are indexed by the augmented set. This technique can improve the performance of endmember extraction; however, the results may vary due to the inherent randomness of the selection process. To mitigate this variability, we employ an averaging strategy: the endmember extraction process is repeated multiple times with different random selections, and the resulting estimates are then averaged.
We summarize the overall procedure of REDIC in Algorithm 3.
Algorithm 3 REDIC: Data-reduced self-dictionary method for endmember extraction
Input: $A\in\mathbb{R}^{d\times n}$; positive integer $r$; parameters $\lambda,\tau$
Output: $\hat{\bm{w}}_{1},\ldots,\hat{\bm{w}}_{r}\in\mathbb{R}^{d}$
  1. 1.
    Compute the top-$r$ truncated SVD $A_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$ of $A$ and construct $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$.
  2. 2.
    Run DRS on $A^{\prime}$ to obtain $\mathcal{K}\in\Gamma(A^{\prime})$.
  3. 3.
    For $j=1,\ldots,\tau$, do:
    1. 3-1.
      Randomly select $\lambda$ elements from $n]\setminus\mathcal{K}$ and construct the set $\mathcal{K}_{\mathrm{add}}$.
    2. 3-2.
      Construct the model $\mathsf{H}$ using $A^{\prime}(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $r$, solve it with RCE, and perform the postprocessing step using method-C to obtain $r$ indices $i_{1},\ldots,i_{r}$.
    3. 3-3.
      Set $W_{j}=[\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}]$ for the columns $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ of $A$.
  4. 4.
    If $\tau=1$, set $\hat{W}=W_{1}$ and go to Step 7; otherwise go to Step 5.
  5. 5.
    Initialize $C\leftarrow W_{1}$. For $j=2,\ldots,\tau$, do:
    1. 5-1.
      Update $C\leftarrow\tfrac{1}{j-1}(W_{1}+\cdots+W_{j-1})$.
    2. 5-2.
      Rearrange the columns of $W_{j}$ so as to minimize the sum of MRSA values between corresponding columns of $C$ and $W_{j}$.
  6. 6.
    Set $\hat{W}=\tfrac{1}{\tau}(W_{1}+\cdots+W_{\tau})$.
  7. 7.
    Return the columns $\hat{\bm{w}}_{1},\ldots,\hat{\bm{w}}_{r}$ of $\hat{W}$ and terminate.
5 Experiments
5.1 Objectives and Experimental Setup
We implemented REDIC in MATLAB and evaluated its performance through numerical experiments. The first experiment (Section 5.2) was conducted to examine the output of the DRS module incorporated into REDIC. In particular, we investigated the following two aspects: the scale of the matrix obtained by applying DRS, and whether this matrix retains the endmembers present in the original HSI data. The second experiment (Section 5.3) was conducted to evaluate the performance of REDIC in terms of both endmember extraction accuracy and computation time. All experiments were carried out in MATLAB on dual Intel Xeon Gold 6336Y processors with 256 GB of memory.
In what follows, we describe the implementation details and datasets used in our experiments. REDIC performs DRS, which involves the parameter $p$ that controls the number of clusters in the $k$-means method. We set $p=30$ in all experiments. DRS calls DR to find a set $\mathcal{K}\in\Gamma(A)$ for a matrix $A$. DR determines the feasibility of whether $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$ by solving problem (5). We used the MATLAB function lsqnonneg, which implements the active set method, to solve this problem. The tolerance parameter $\epsilon_{\mathrm{feas}}$ in the feasibility test was set to $10^{-8}$.
REDIC also performs RCE to solve the model $\mathsf{H}$. For the implementation of RCE, we used a reorganized and speed-improved version of the MATLAB code developed in [25], in which CPLEX is employed as the LP solver. The comparison of computation time with the previous version of the RCE code is described in Remark 1. RCE first solves a small-scale subproblem of $\mathsf{H}$ and then gradually increases the problem size until the optimal solution of $\mathsf{H}$ is obtained. It includes the parameters $\zeta$ and $\eta$ to control the size of the initial subproblem. These parameter values were set as follows to maintain consistency with the experimental setup in [25]: for an input matrix with $n$ columns, if $n\leq 300$, then $(\zeta,\eta)=(0,n)$; if $300<n\leq 50000$, then $(\zeta,\eta)=(10,100)$; and if $n>50000$, then $(\zeta,\eta)=(50,300)$.
We employed three real HSI datasets in our experiments: Jasper Ridge, Samson, and Urban, which are widely used for evaluating endmember extraction methods. These datasets were constructed following the procedure described in [29] and are the same as those used in [25]. The details of each dataset are as follows:
  • Jasper Ridge: The Jasper Ridge HSI was acquired by the AVIRIS sensor and consists of $512\times 614$ pixels with 224 bands covering wavelengths from 360 nm to 2500 nm. We used a subimage of size $100\times 100$ pixels extracted from the original image, containing 198 valid bands after removing the degraded ones. This region includes four endmembers: Tree, Soil, Water, and Road.
  • Samson: The Samson HSI consists of $952\times 952$ pixels with 156 bands covering wavelengths from 400 nm to 900 nm. We used a subimage of size $95\times 95$ pixels from the original image with all 156 bands retained. This region includes three endmembers: Soil, Tree, and Water.
  • Urban: The Urban HSI was acquired by the HYDICE sensor and consists of $307\times 307$ pixels with 210 bands covering wavelengths from 400 nm to 2500 nm. We used the entire image with 162 valid bands after removing the degraded ones. This region contains four to six endmembers. Following the setting in [15, 25], we set the number of endmembers to six: Asphalt, Grass, Tree, Roof 1, Roof 2, and Soil.
Table 1 summarizes, for each dataset, the size of the HSI matrix $A\in\mathbb{R}^{d\times n}$, the number of endmembers $r$, and the specific names of the endmembers considered in our experiments. Figure 2 shows the RGB images of these datasets.
Table 1: Specifications of the HSI datasets.
Dataset $(d,n,r)$ Endmembers
Jasper Ridge $(198,10000,4)$ Tree, Soil, Road
Samson $(156,9025,3)$ Soil, Tree, Water
Urban $(162,94249,6)$ Asphalt, Grass, Tree, Roof 1, Roof 2, Soil
Figure 2: RGB images of the Jasper Ridge (left), Samson (center), and Urban (right) datasets.
5.2 Performance Evaluation of DRS
First, we evaluated the performance of the DRS module incorporated into REDIC using synthetic datasets. The datasets were derived from the three real HSI datasets, as described in Section 5.1, following the procedure presented in Section VII-B of [25]. To make this paper self-contained, we summarize the procedure here. The endmember signatures for each HSI dataset were identified in [29], and we used them to generate the synthetic datasets. Let $A^{\mathrm{real}}\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given HSI dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ represent the identified endmember signatures. The following procedure was used to generate $W\in\mathbb{R}^{d\times r}$, $H\in\mathbb{R}^{r\times n}$, and $V\in\mathbb{R}^{d\times n}$:
  1. 1.
    Normalize all columns of $A^{\mathrm{real}}$ to have unit $L_{1}$ norm.
  2. 2.
    For $j=1,\ldots,r$, compute
    $i_{j}=\arg\min_{i\in[n]}\operatorname{MRSA}(\bm{w}_{j}^{\mathrm{ident}},\bm{a}^{\mathrm{real}}_{i}),$
    where $\bm{a}^{\mathrm{real}}_{i}$ is the $i$th column of $A^{\mathrm{real}}$.
  3. 3.
    Set $W=A^{\mathrm{real}}(\{i_{1},\ldots,i_{r}\})$.
  4. 4.
    Compute the optimal solution $X$ to the following convex optimization problem:
    $\min_{X\in\mathbb{R}^{r\times n}}\ \|A^{\mathrm{real}}-WX\|_{F}^{2}\quad\text{subject to}\quad\bm{1}^{\top}X=\bm{1}^{\top},\ X\geq O,$
    and set $X(\{i_{1},\ldots,i_{r}\})=I$.
  5. 5.
    Set $H=X$ and $V=A^{\mathrm{real}}-WH$.
Three synthetic datasets were constructed using the matrices $W$, $H$, and $V$ obtained through the above procedure: Dataset 1 from Jasper Ridge, Dataset 2 from Samson, and Dataset 3 from Urban. Each dataset contained HSI matrices generated as $A=WH+(\nu/\|V\|_{1})\cdot V\in\mathbb{R}^{d\times n},$ where the noise intensity level $\nu$ was varied from 0 to $1.5$ in increments of $0.1$, resulting in 16 HSI matrices with different noise levels. The matrix size $(d,n)$ and the number of endmembers $r$ for Datasets 1–3 are identical to those of the corresponding original HSI datasets. When $\nu=\|V\|_{1}$, the HSI matrix $A$ coincides with $A^{\mathrm{real}}$, whose columns are all normalized to have unit $L_{1}$ norm. In particular, $\|V\|_{1}$ equals 0.61, 0.15, and 1.15 for Datasets 1–3, respectively.
We examined the output of DRS during the execution of REDIC. To avoid ambiguity, we briefly describe the tested algorithm below. The algorithm takes as input a matrix $A$ and a parameter $\lambda$:
  1. 1.
    Perform Steps 1 and 2 of REDIC to obtain $\mathcal{K}\in\Gamma(A^{\prime})$, where $A^{\prime}$ is the reduced matrix obtained through the SVD-based dimensionality reduction in Step 1.
  2. 2.
    If $\lambda=0$, then set $\mathcal{K}_{\mathrm{add}}=\emptyset$; otherwise, repeat the process 50 times to generate 50 sets of $\mathcal{K}_{\mathrm{add}}$, each formed by randomly selecting $\lambda$ elements from $n]\setminus\mathcal{K}$.
After completing the above procedure, we evaluated (i) the size of $\mathcal{K}$, (ii) the reconstruction error of $A^{\prime}$ based on $A^{\prime}(\mathcal{K})$, and (iii) the distance between $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $W$.
The reconstruction error was introduced to assess how well the original reduced matrix $A^{\prime}$ can be reconstructed from the submatrix $A^{\prime}(\mathcal{K})$. The metric is defined as follows. Recall that $A^{\prime}$ is of size $r\times n$, and let its columns be denoted by $\bm{a}^{\prime}_{1},\ldots,\bm{a}^{\prime}_{n}$. First, we compute the optimal solutions $\bm{x}_{1},\ldots,\bm{x}_{n}$ to the following nonnegative least-squares problems:
$\bm{x}_{i}=\arg\min_{\bm{x}\geq\bm{0}}\,\|A^{\prime}(\mathcal{K})\bm{x}-\bm{a}^{\prime}_{i}\|_{2}^{2}\quad\text{for }i\in[n].$
The reconstruction error of $A^{\prime}$ based on $A^{\prime}(\mathcal{K})$ is then defined as
$\sqrt{\frac{1}{rn}\sum_{i=1}^{n}\|A^{\prime}(\mathcal{K})\bm{x}_{i}-\bm{a}^{\prime}_{i}\|_{2}^{2}}.$
The distance between $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $W$ was introduced to quantify how closely each column of $W$ can be matched by at least one column of $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$. We define this metric as follows. To simplify the notation, let $\operatorname{dist}(\bm{a},\bm{b})$ denote either the $L_{1}$ distance or the MRSA value between $\bm{a},\bm{b}\in\mathbb{R}^{d}$. For each column $\bm{w}_{i}$ of $W$ with $i\in[r]$, we select an index $k_{i}$ from $\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}$ corresponding to the column of $A$ that is closest to $\bm{w}_{i}$ under the chosen metric:
$k_{i}=\arg\min_{k\in\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}}\,\operatorname{dist}(\bm{w}_{i},\bm{a}_{k})\quad\text{for }i\in[r].$
We then define the distance between $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $W$ as the mean of these distances,
$\frac{1}{r}\sum_{i=1}^{r}\operatorname{dist}(\bm{w}_{i},\bm{a}_{k_{i}}),$
which we refer to as the $L_{1}$ distance (resp. MRSA) of the DRS output $\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}$ with respect to the reference endmember matrix $W$, when $\operatorname{dist}$ represents the $L_{1}$ distance (resp. MRSA value). These two metrics were used to assess the similarity between $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $W$.
The above algorithm was executed by varying $\lambda$ from 0 to 250 in increments of 50 for Datasets 1 and 2, and from 0 to 2500 in increments of 500 for Dataset 3. Each value of $\lambda$ corresponds to approximately $0\%$$2.5\%$ of the number of columns $n$ in each dataset, with increments of 0.5%. In addition, for Dataset 3, $\lambda$ was also varied from 0 to 12500 in increments of 2500, corresponding to approximately $0\%$$12.5\%$ of the number of columns $n$, with increments of 2.5%.
(a) Dataset 1
(b) Dataset 2
Figure 3: Experimental results for Datasets 1 and 2. Each pair of plots shows the $L_{1}$ distance (left) and the MRSA distance (right) of the DRS output $\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}$. For $\lambda>0$, each plot shows the mean of the corresponding distance over 50 realizations of $\mathcal{K}_{\mathrm{add}}$.
(a) Dataset 3 with $\lambda$ ranging from 0 to 2500 in increments of 500
(b) Dataset 3 with $\lambda$ ranging from 0 to 10000 in increments of 2500
Figure 4: Experimental results for Dataset 3 with $\lambda$ ranging from 0 to 2500 in increments of 500 (top) and from 0 to 10000 in increments of 2500 (bottom). Each pair of plots shows the $L_{1}$ distance (left) and the MRSA distance (right) of the DRS output $\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}$. For $\lambda>0$, each plot shows the mean of the corresponding distance over 50 realizations of $\mathcal{K}_{\mathrm{add}}$.
We first report the size of the output set $\mathcal{K}$ obtained by DRS. For each dataset, this size coincided with the number of endmembers when $\nu=0$. When $\nu>0$, the average sizes of $\mathcal{K}$ were 64.6 for Dataset 1, 16.4 for Dataset 2, and 546.8 for Dataset 3. In other words, DRS retained only about 0.2%–0.6% of the columns for each dataset. Next, we report the reconstruction error of $A^{\prime}$ based on $A^{\prime}(\mathcal{K})$. The error values were smaller than $10^{-8}$ for all datasets and all noise intensity levels $\nu$. We then report the distance between $A(\mathcal{K}\cup\mathcal{K}_{\mathrm{add}})$ and $W$. Figures 3 and 4 summarize the experimental results for Datasets 1–3, showing the $L_{1}$ and MRSA distances of the DRS output $\mathcal{K}\cup\mathcal{K}_{\mathrm{add}}$ at each noise intensity level $\nu$. For $\lambda>0$, each plot shows the mean of the corresponding distance over 50 realizations of $\mathcal{K}_{\mathrm{add}}$ generated for each value of $\lambda$. It can be observed from these figures that the distances decrease as $\lambda$ increases for all datasets. In particular, even when $\lambda$ is set to 1% of $n$, the distances are substantially reduced compared with those at $\lambda=0$. Finally, we report the computational time required to execute DRS. Specifically, we measured the elapsed time for Steps 1 and 2 of REDIC. The average elapsed times over the 16 noise intensity levels $\nu$ were 1–2 seconds for Datasets 1 and 2, and 2 minutes for Dataset 3.
5.3 Performance Evaluation of REDIC
Next, we evaluated the performance of REDIC using the three real HSI datasets described in Section 5.1. The following metric was employed to assess the accuracy of endmember extraction. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}\in\mathbb{R}^{d}$ represent the endmember signatures identified in [29]. We compute the column indices $i_{1},\ldots,i_{r}$ of $A$ corresponding to $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ under the MRSA metric as
$i_{j}=\arg\min_{i\in[n]}\operatorname{MRSA}(\bm{w}_{j}^{\mathrm{ident}},\bm{a}_{i})\quad\text{for }j\in[r],$
and regard $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ as the reference endmember signatures. This procedure was introduced to ensure the reproducibility of the experimental results. Let $\bm{w}_{1},\ldots,\bm{w}_{r}$ denote the reference endmember signatures, and let $\hat{\bm{w}}_{1},\ldots,\hat{\bm{w}}_{r}$ denote the estimated endmember signatures obtained by a given endmember extraction method. We then determine a permutation $\sigma$ on the set $[r]$ according to
$\sigma=\arg\min_{\sigma\in\mathcal{S}_{r}}\sum_{j=1}^{r}\operatorname{MRSA}(\bm{w}_{\sigma(j)},\hat{\bm{w}}_{j}),$
where $\mathcal{S}_{r}$ denotes the set of all permutations on $[r]$. Such a permutation can be obtained by solving an assignment problem. The MRSA values between the reference and estimated endmembers, $\operatorname{MRSA}(\bm{w}_{\sigma(j)},\hat{\bm{w}}_{j})$ for $j\in[r]$, and their mean value,
$\frac{1}{r}\sum_{j=1}^{r}\operatorname{MRSA}(\bm{w}_{\sigma(j)},\hat{\bm{w}}_{j}),$
are then computed. We define this mean value as the MRSA score of the method.
This experiment was based on the EEHT-C experiment on the Urban dataset presented in [25], where a data-specific preprocessing technique was applied prior to executing EEHT-C to enhance its performance (see Section VII-C of [25] for details). The preprocessing technique involves two parameters, $\phi$ and $\omega$. In this experiment, before executing endmember extraction, we applied the same preprocessing to the Urban dataset, using $(\phi,\omega)=(0.4,0.1)$ as in [25]. This procedure eliminated 2995 columns from the original HSI matrix, which contained 94249 columns, resulting in a total of 91254 columns. For Jasper Ridge and Samson, no such preprocessing was applied.
Before presenting the experimental results for REDIC, we first show the specifications of the DRS output $\mathcal{K}$ for each dataset in Table 2. These results were obtained by executing Steps 1 and 2 of REDIC with $\lambda=0$. It can be observed that only about 0.2%–0.5% of the columns were retained for each dataset, that the reduced matrix $A^{\prime}$ can be reconstructed from the submatrix $A^{\prime}(\mathcal{K})$ with negligible error, and that the reduced matrices contain columns sufficiently similar to the reference endmember signatures. It should be noted that these MRSA distances serve as lower bounds on the MRSA scores achievable by applying endmember extraction methods to the reduced matrices. Figure 1, shown in Section 1, was generated using the DRS output $\mathcal{K}$ produced by executing Steps 1 and 2 of REDIC with $\lambda=0$ for the Samson dataset.
Table 2: Specifications of the DRS output $\mathcal{K}$ for each dataset.
Jasper Ridge Samson Urban
Number of elements 53 20 483
MRSA distance 5.96 2.48 6.15
Reconstruction error $2.11\times 10^{-12}$ $2.52\times 10^{-13}$ $1.05\times 10^{-16}$
Table 3: Experimental results for the three datasets: MRSA scores and elapsed times. The first row for each dataset shows the results of the LP method, and the subsequent rows show those of REDIC with various $(\lambda,\tau)$ settings. For REDIC with $\lambda>0$, the mean and standard deviation over 50 runs are shown.
(a) Jasper Ridge
MRSA score Elapsed time (s) LP method $6.82$ $271.9$ ($\lambda,\tau$) = (0, 1) $12.24$ $6.3$ ($\lambda,\tau$) = (50, 1) $12.24\pm 0.00$.00) $4.9\pm 0.0$.4) ($\lambda,\tau$) = (50, 5) $12.24\pm 0.00$.00) $5.3\pm 0.0$.3) ($\lambda,\tau$) = (100, 1) $8.49\pm 0.01$.10) $5.5\pm 0.0$.4) ($\lambda,\tau$) = (100, 5) $7.10\pm 0.00$.87) $8.2\pm 0.0$.4) ($\lambda,\tau$) = (150, 1) $7.39\pm 0.01$.42) $7.5\pm 0.0$.4) ($\lambda,\tau$) = (150, 5) $6.11\pm 0.00$.63) $18.8\pm 0.0$.7) ($\lambda,\tau$) = (200, 1) $7.04\pm 0.01$.22) $10.2\pm 0.0$.5) ($\lambda,\tau$) = (200, 5) $5.77\pm 0.00$.48) $32.7\pm 0.1$.0) ($\lambda,\tau$) = (250, 1) $6.88\pm 0.01$.25) $14.3\pm 0.0$.8) ($\lambda,\tau$) = (250, 5) $5.68\pm 0.00$.56) $52.3\pm 0.1$.7)
(b) Samson
MRSA score Elapsed time (s) LP method $3.34$ $31.2$ ($\lambda,\tau$) = (0, 1) $6.14$ $1.6$ ($\lambda,\tau$) = (50, 1) $3.89\pm 0.00$.99) $1.5\pm 0.0$.1) ($\lambda,\tau$) = (50, 5) $3.72\pm 0.00$.48) $1.7\pm 0.0$.0) ($\lambda,\tau$) = (100, 1) $3.23\pm 0.00$.59) $2.2\pm 0.0$.1) ($\lambda,\tau$) = (100, 5) $3.05\pm 0.00$.28) $5.9\pm 0.0$.5) ($\lambda,\tau$) = (150, 1) $2.99\pm 0.00$.43) $4.2\pm 0.0$.3) ($\lambda,\tau$) = (150, 5) $2.90\pm 0.00$.25) $15.7\pm 0.0$.8) ($\lambda,\tau$) = (200, 1) $2.91\pm 0.00$.34) $7.0\pm 0.0$.6) ($\lambda,\tau$) = (200, 5) $77\pm 0.00.22$) $29.5\pm 0.1$.1) ($\lambda,\tau$) = (250, 1) $2.87\pm 0.00$.37) $11.2\pm 0.1$.0) ($\lambda,\tau$) = (250, 5) $68\pm 0.00.22$) $52.6\pm 0.2$.4)
(c) Urban
MRSA score Elapsed time (s) LP method $7.90$ $29395.5$ ($\lambda,\tau$) = (0, 1) $15.86$ $126.5$ ($\lambda,\tau$) = (500, 1) $16.06\pm 0.00$.78) $112.7\pm 0.3$.0) ($\lambda,\tau$) = (500, 5) $15.41\pm 0.00$.45) $141.1\pm 0.5$.3) ($\lambda,\tau$) = (1000, 1) $13.55\pm 0.01$.71) $113.3\pm 0.3$.7) ($\lambda,\tau$) = (1000, 5) $12.25\pm 0.01$.03) $164.3\pm 0.7$.1) ($\lambda,\tau$) = (1500, 1) $12.43\pm 0.02$.06) $116.8\pm 0.6$.3) ($\lambda,\tau$) = (1500, 5) $10.91\pm 0.01$.05) $179.0\pm 1.2$.7) ($\lambda,\tau$) = (2000, 1) $11.41\pm 0.01$.91) $123.7\pm 0.5$.0) ($\lambda,\tau$) = (2000, 5) $9.93\pm 0.00$.96) $210.8\pm 1.2$.4) ($\lambda,\tau$) = (2500, 1) $10.59\pm 0.01$.64) $122.0\pm 0.5$.0) ($\lambda,\tau$) = (2500, 5) $9.29\pm 0.00$.97) $209.1\pm 1.0$.5) ($\lambda,\tau$) = (5000, 1) $9.74\pm 0.01$.66) $162.9\pm 2.6$.9) ($\lambda,\tau$) = (5000, 5) $8.41\pm 0.00$.59) $435.0\pm 7.3$.9) ($\lambda,\tau$) = (7500, 1) $9.43\pm 0.01$.04) $218.6\pm 7.4$.1) ($\lambda,\tau$) = (7500, 5) $8.16\pm 0.00$.48) $765.7\pm 24.4$.4) ($\lambda,\tau$) = (10000, 1) $9.24\pm 0.00$.91) $275.8\pm 9.9$.8) ($\lambda,\tau$) = (10000, 5) $7.98\pm 0.00$.35) $1009.0\pm 27.0$.7) ($\lambda,\tau$) = (12500, 1) $9.14\pm 0.00$.73) $361.6\pm 17.5$.5) ($\lambda,\tau$) = (12500, 5) $7.99\pm 0.00$.39) $1339.5\pm 41.1$.6)
We then present the experimental results for REDIC. The REDIC and LP methods were applied to each HSI dataset, and their MRSA scores and computational times were compared. The LP method examined here is identical to EEHT-C proposed in [25], whose procedure was reviewed in Section 4. The parameters $\lambda$ and $\tau$ in REDIC were set as follows: $\tau$ took values in $\{1,5\}$, and $\lambda$ was varied in the same manner as in the first experiment described in Section 5.2. When $\lambda>0$, REDIC was executed 50 times for each of the three HSI datasets, and the mean and standard deviation of the MRSA scores and elapsed times across the 50 runs were calculated. Table 3 summarizes the experimental results. We can make the following observations from these results:
  • For all datasets, the MRSA scores of REDIC with $\lambda=0$ are higher than those of the LP method, and noticeable gaps exist between these MRSA scores and the MRSA distances of the DRS output reported in Table 2.
  • The MRSA scores of REDIC tend to decrease as $\lambda$ increases for all datasets.
  • When $\lambda>0$, the mean and standard deviation of the MRSA scores with $\tau=5$ are generally lower than those with $\tau=1$.
  • For Jasper Ridge and Samson, REDIC with $(\lambda,\tau)=(100,5)$ achieves nearly the same MRSA scores as the LP method, while reducing the computation time and achieving speed-ups of 33 and 5 times, respectively. In other words, adding only about 1% of the columns was sufficient to extract endmembers with accuracy comparable to that of the LP method. Moreover, the MRSA scores decrease further as $\lambda$ increases.
  • For Urban, REDIC with $(\lambda,\tau)=(10000,5)$ achieves nearly the same MRSA scores as the LP method, with a speed-up of 29 times. In contrast to Jasper Ridge and Samson, about 10% of the columns had to be added to achieve comparable accuracy.
The LP method took 8 hours to process the Urban dataset, whereas REDIC completed the same task in only 16 minutes, without compromising endmember extraction accuracy. These results suggest that incorporating DRS into self-dictionary methods enables efficient processing of large-scale datasets within a reasonable time frame.
Remark 1.
The previous version of the MATLAB code for the LP method (i.e., EEHT-C) developed in [25] required 10 hours to process the Urban dataset. In contrast, the reorganized and speed-improved version of the code used in this experiment completed the same task in 8 hours.
6 Analysis of Reduced Hyperspectral Image Data
Throughout this section, we assume that $A\in\mathbb{R}^{d\times n}$ is a nearly $r$-separable matrix of the form $A=M+V$, where $M\in\mathbb{R}^{d\times n}_{+}$ is $r$-separable and can be written as $M=WH$, as given in (4), and $V\in\mathbb{R}^{d\times n}$ is noise. Before presenting the proof of Theorem 1, we establish some properties of the parameter $\rho$.
Lemma 1.
The following statements hold for $Z\in\mathbb{R}^{m\times n}$:
  1. (a)
    $\rho(Z)=0$ if and only if the $n$ columns of $Z$ are linearly dependent.
  2. (b)
    $\rho(Z)>0$ if and only if the $n$ columns of $Z$ are linearly independent.
Proof.
Since part (b) is immediate from part (a), we prove only part (a). We first show the “only if” direction. There exists $\bm{x}\in\mathbb{R}^{n}$ such that $Z\bm{x}=\bm{0}$ and $\|\bm{x}\|_{1}=1$, which implies that the $n$ columns of $Z$ are linearly dependent. We next prove the “if” direction. There exists $\bm{x}\neq\bm{0}$ such that $Z\bm{x}=\bm{0}$. Let $\bar{\bm{x}}=\bm{x}/\|\bm{x}\|_{1}$, Then $Z\bar{\bm{x}}=\bm{0}$ and $\|\bar{\bm{x}}\|_{1}=1$. Accordingly, $\rho(Z)=0$. ∎
Lemma 2.
Let $Z\in\mathbb{R}^{m\times n}$ satisfy $\|Z\|_{1}=1$. Then $\rho(Z)\leq 1$ holds.
Proof.
For any $\bm{x}\in\mathbb{R}^{n}$ such that $\|\bm{x}\|_{1}=1$, we have $\|Z\bm{x}\|_{1}\leq\|Z\|_{1}\|\bm{x}\|_{1}=1$. Accordingly, $\rho(Z)=\min_{\|\bm{x}\|_{1}=1}\|Z\bm{x}\|_{1}\leq 1$. ∎
As an immediate consequence of Lemma 1, we obtain the following result:
Corollary 1.
If $\rho(Z)>0$ for $Z\in\mathbb{R}^{m\times n}$, then $m\geq n$.
Proof.
If $\rho(Z)>0$, then by part (b) of Lemma 1, the $n$ columns of $Z$ are linearly independent, which implies that $\operatorname{rank}(Z)=n$. Since $\operatorname{rank}(Z)=\min\{m,n\}\leq m$, it follows that $m\geq n$. ∎
6.1 Proof of Theorem 1
To prove Theorem 1, we establish two propositions: Propositions 1 and 2. In the first proposition, we focus on the factor $H$ of the $r$-separable matrix $M=WH$ contained in $A$. We choose $\mathcal{K}$ from $\Gamma(A)$ and consider, for each $j$, the maximum value in the $j$th row of $H(\mathcal{K})$ and the index at which this maximum is attained. We denote these respectively by $\mu(j)$ and $k(j)$. More precisely, given $\mathcal{K}\in\Gamma(A)$, define
$\mu(j)$ $=\max_{k\in\mathcal{K}}H(j,k),$ (6)
$k(j)$ $=\arg\max_{k\in\mathcal{K}}H(j,k)$ (7)
for each $j\in[r]$.
In the first proposition, we show that if $\mu(j)$ exceeds $1/2$, then the indices $k(j)$ are all distinct, and the $L_{1}$ distance between $\bm{w}_{j}$ and $\bm{a}_{k(j)}$ is bounded above by $2(1-\mu(j))$ plus the noise level $\epsilon$ involved in $A$. This proposition provides a framework for proving Theorem 1.
Proposition 1.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. Let $\mathcal{K}\in\Gamma(A)$. If $\epsilon<\rho(W)$ and $\mu(j)>1/2$ for each $j\in[r]$, then
  • the elements $k(1),\ldots,k(r)$ of $\mathcal{K}$ are all distinct, and
  • $\|\bm{w}_{j}-\bm{a}_{k(j)}\|_{1}\leq 2(1-\mu(j))+\epsilon$ for each $j\in[r]$.
The proof is given in Section 6.2. Next, we show in the second proposition that, if $\epsilon$ is below a certain threshold, then $\mu(j)$ exceeds $1/2$.
Proposition 2.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. Let $\mathcal{K}\in\Gamma(A)$. If $\epsilon<\rho(W)/9$, then
$1-\mu(j)\leq\frac{4\epsilon}{\rho(W)(1-\epsilon)}<\frac{1}{2}$
for each $j\in[r]$.
The proof is given in Section 6.2. We now prove Theorem 1 by combining Propositions 1 and 2.
(Proof of Theorem 1).
It follows from Propositions 1 and 2 that $\mathcal{K}\in\Gamma(A)$ contains $r$ distinct indices $k_{1},\ldots,k_{r}$, which satisfy
$\|\bm{w}_{j}-\bm{a}_{k_{j}}\|_{1}\leq 2(1-\mu(j))+\epsilon\leq\frac{8}{\rho(W)}\frac{\epsilon}{1-\epsilon}+\epsilon$
for each $j\in[r]$. Lemma 2, together with Assumption 1(a), implies that $\rho(W)\leq 1$. Therefore, $0\leq\epsilon<\rho(W)/9\leq 1/9$. It then follows that $\epsilon/(1-\epsilon)<9\epsilon/8$. Consequently, we obtain $\|\bm{w}_{j}-\bm{a}_{k_{j}}\|_{1}<(9/\rho(W)+1)\epsilon$ for each $j\in[r]$. ∎
6.2 Proof of Proposition 1
We begin by evaluating the $L_{1}$ distance between $\bm{w}_{j}$ and $\bm{a}_{i}$. We first show that this distance can be upper-bounded using the $(j,i)$th element of $H$ and the noise level $\epsilon$ under Assumption 1.
Lemma 3.
Let the nearly $r$-separable matrix $A=M+V$, with $M=WH$, satisfy Assumption 1. Then $\|\bm{w}_{j}-\bm{a}_{i}\|_{1}\leq 2(1-H(j,i))+\epsilon$ holds for each $i\in[n]$ and each $j\in[r]$.
Proof.
The $i$th column $\bm{a}_{i}$ of $A$ can be expressed as $\bm{a}_{i}=\sum_{u=1}^{r}H(u,i)\bm{w}_{u}+\bm{v}_{i}$. In addition, by $H\geq O$ and Assumption 1(a), we have
$\sum_{u=1}^{r}H(u,i)=1\operatorname{\Leftrightarrow}1-H(j,i)=\sum_{u\neq j}H(u,i)\geq 0.$
Accordingly, we can bound $\|\bm{w}_{j}-\bm{a}_{i}\|_{1}$ from above as follows:
$\|\bm{w}_{j}-\bm{a}_{i}\|_{1}$ $=\|\bm{w}_{j}-\sum_{u=1}^{r}H(u,i)\bm{w}_{u}-\bm{v}_{i}\|_{1}$
$=\|(1-H(j,i))\bm{w}_{j}-\sum_{u\neq j}H(u,i)\bm{w}_{u}-\bm{v}_{i}\|_{1}$
$\leq(1-H(j,i))\|\bm{w}_{j}\|_{1}+\sum_{u\neq j}H(u,i)\|\bm{w}_{u}\|_{1}+\|\bm{v}_{i}\|_{1}$
$\leq 1-H(j,i)+\sum_{u\neq j}H(u,i)+\epsilon$
$=2(1-H(j,i))+\epsilon$
where the first inequality follows from $H\geq O$ and $1-H(j,i)\geq 0$ as shown above, and the second inequality follows from Assumption 1. ∎
Let $\mathcal{K}\in\Gamma(A)$. This lemma implies that, for each $k\in\mathcal{K}$, the $L_{1}$ distance between $\bm{w}_{j}$ and $\bm{a}_{k}$ satisfies $\|\bm{w}_{j}-\bm{a}_{k}\|_{1}\leq 2(1-H(j,k))+\epsilon$. In particular, this upper bound is minimized when $k=k(j)$, where $k(j)$ is defined as in (7). Hence, the column $\bm{a}_{k(j)}$ is located near $\bm{w}_{j}$, and the $L_{1}$ distance between them is bounded as
$\|\bm{w}_{j}-\bm{a}_{k(j)}\|_{1}\leq 2(1-\mu(j))+\epsilon$
for each $j\in[r]$.
We next show that if $\mu(j)>1/2$ for all $j\in[r]$, then the indices $k(1),\ldots,k(r)$ are all distinct. In what follows, we regard $k(j)$ as a map from $[r]$ to $\mathcal{K}$. We show in Lemma 4 that any $\mathcal{K}\in\Gamma(A)$ contains at least $r$ elements. Hence, it is possible for an injective map from $[r]$ to $\mathcal{K}$ to exist. We then show in Lemma 5 that if $\mu(j)>1/2$ for all $j\in[r]$, then the map $k:[r]\rightarrow\mathcal{K}$ is in fact injective.
Lemma 4.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. If $\epsilon<\rho(W)$, then any $\mathcal{K}\in\Gamma(A)$ satisfies $|\mathcal{K}|\geq r$.
Proof.
Since $\Gamma(A)\neq\emptyset$, there exists $\mathcal{K}_{*}\in\Gamma(A)$ such that $|\mathcal{K}_{*}|\leq|\mathcal{K}|$ for any $\mathcal{K}\in\Gamma(A)$. Hence, it suffices to show that $|\mathcal{K}_{*}|\geq r$. We prove this inequality by contradiction. Suppose, contrary to the claim, that $|\mathcal{K}_{*}|<r$. There are the columns $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ of $A$ such that $\bm{a}_{i_{j}}=\bm{w}_{j}+\bm{v}_{i_{j}}$ for $j\in[r]$. By letting $\mathcal{I}=\{i_{1},\ldots,i_{r}\}$, we express these columns in a matrix form as $A(\mathcal{I})=W+V(\mathcal{I})$. For any $\bm{x}\in\mathbb{R}^{r}$ with $\|\bm{x}\|_{1}=1$,
$\|A(\mathcal{I})\bm{x}\|_{1}=\|W\bm{x}+V(\mathcal{I})\bm{x}\|_{1}\geq\|W\bm{x}\|_{1}-\|V(\mathcal{I})\bm{x}\|_{1}\geq\rho(W)-\epsilon.$
The last inequality follows from the definition of $\rho(W)$ and Assumption 1(b). We thus get
$\rho(A(\mathcal{I}))=\min_{\|\bm{x}\|_{1}=1}\|A(\mathcal{I})\bm{x}\|_{1}\geq\rho(W)-\epsilon>0.$
Taking into account Lemma 1(b), we obtain $\operatorname{rank}(A(\mathcal{I}))=r$. Meanwhile, since $\bm{a}_{1},\ldots,\bm{a}_{n}\in\operatorname{cone}(A)=\operatorname{cone}A((\mathcal{K}_{*}))$, we can express $A(\mathcal{I})$ as $A(\mathcal{I})=A(\mathcal{K}_{*})X$ using $X\in\mathbb{R}^{|\mathcal{K}_{*}|\times r}_{+}$, which implies that
$\operatorname{rank}(A(\mathcal{I}))=\operatorname{rank}(A(\mathcal{K}_{*})X)\leq\operatorname{rank}(A(\mathcal{K}_{*}))\leq\min\{d,|\mathcal{K}_{*}|\}=|\mathcal{K}_{*}|.$
The last equality follows from the assumption $|\mathcal{K}_{*}|<r$ and from $r\leq d$, which holds by Corollary 1 together with $0\leq\epsilon<\rho(W)$. Thus we obtain the contradiction
$\operatorname{rank}(A(\mathcal{I}))=r>|\mathcal{K}_{*}|\geq\operatorname{rank}(A(\mathcal{I})).$
Therefore, $|\mathcal{K}_{*}|\geq r$ must hold. ∎
Lemma 5.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. Let $\mathcal{K}\in\Gamma(A)$. If $\epsilon<\rho(W)$ and $\mu(j)>1/2$ for each $j\in[r]$, then the map $k:[r]\rightarrow\mathcal{K}$ is injective.
Proof.
Since $\epsilon<\rho(W)$, Lemma 4 implies that an injective map from $[r]$ to $\mathcal{K}$ can exist. We prove by contradiction that if $j\neq j^{\prime}$, then $k(j)\neq k(j^{\prime})$. Suppose, contrary to the claim, that $k=k(j)=k(j^{\prime})$ for some $j\neq j^{\prime}$. Since $\mu(j)>1/2$ for each $j\in[r]$, this assumption implies that
$H(j,k)=H(j,k(j))=\mu(j)>1/2,\qquad H(j^{\prime},k)=H(j^{\prime},k(j^{\prime}))=\mu(j^{\prime})>1/2.$
Accordingly, since $H\geq O$, we obtain $\sum_{u=1}^{r}H(u,k)>1$. On the other hand, by Assumption 1(a) and $H\geq O$, we have $\sum_{u=1}^{r}H(u,k)=1$. This yields the contradiction
$1=\sum_{u=1}^{r}H(u,k)>1.$
Therefore, the assumption is false, and hence $k(j)\neq k(j^{\prime})$ whenever $j\neq j^{\prime}$. ∎
We now prove Proposition 1.
(Proof of Proposition 1).
The result follows from Lemmas 3 and 5. ∎
6.3 Proof of Proposition 2
We estimate an upper bound on $1-\mu(j)$. The nearly $r$-separable matrix $A$ contains columns $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ such that $\bm{a}_{i_{j}}=\bm{w}_{j}+\bm{v}_{i_{j}}$ for $j\in[r]$. Let $\mathcal{K}\in\Gamma(A)$. Since $\bm{a}_{i_{j}}=\bm{w}_{j}+\bm{v}_{i_{j}}\in\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}))$, there exists some $\bm{c}_{j}\in\mathbb{R}^{|\mathcal{K}|}$ such that
$\bm{w}_{j}+\bm{v}_{i_{j}}=A(\mathcal{K})\bm{c}_{j}\quad\text{and}\quad\bm{c}_{j}\geq\bm{0}$ (8)
for each $j\in[r]$. Using this representation, we derive an upper bound on $1-\mu(j)$ in Lemmas 6 and 7.
Lemma 6.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. Let $\mathcal{K}\in\Gamma(A)$. If $\rho(W)>0$, then for each $j\in[r]$, the following inequality holds:
$1-\mu(j)\leq\frac{2\epsilon+(1+\epsilon)\left|\|\bm{c}_{j}\|_{1}-1\right|}{\rho(W)},$
where $\bm{c}_{j}$ is defined as in (8).
Proof.
Following the notation stated in Section 1, we write $\bm{a}(i)$ for the $i$th entry of a vector $\bm{a}$ in this proof. Let $\mathcal{K}$ contain $s$ elements $k_{1},\ldots,k_{s}$. For $\bm{c}_{j}$, let $\bar{\bm{c}}_{j}=\bm{c}_{j}/\|\bm{c}_{j}\|_{1}$. We rewrite
$\bm{w}_{j}+\bm{v}_{i_{j}}=A(\mathcal{K})\bar{\bm{c}}_{j}+A(\mathcal{K})(\bm{c}_{j}-\bar{\bm{c}}_{j}).$
Each column $\bm{a}_{k_{u}}$ of $A(\mathcal{K})$ can be expressed as $\bm{a}_{k_{u}}=W\bm{h}_{k_{u}}+v_{k_{u}}$ for $u\in[s]$. Hence,
$A(\mathcal{K})\bar{\bm{c}}_{j}=\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{a}_{k_{u}}=\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)(W\bm{h}_{k_{u}}+\bm{v}_{k_{u}})=W\bm{p}+\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{v}_{k_{u}}.$
where $\bm{p}=\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{h}_{k_{u}}\in\mathbb{R}^{r}$. Thus,
$\bm{w}_{j}+\bm{v}_{i_{j}}=W\bm{p}+\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{v}_{k_{u}}+A(\mathcal{K})(\bm{c}_{j}-\bar{\bm{c}}_{j})\operatorname{\Leftrightarrow}\bm{y}=W\bm{q}$
where $\bm{y}=\bm{v}_{i_{j}}-\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{v}_{k_{u}}-A(\mathcal{K})(\bm{c}_{j}-\bar{\bm{c}}_{j})\in\mathbb{R}^{d}$ and $\bm{q}=\bm{p}-\bm{e}_{j}\in\mathbb{R}^{r}$ where $\bm{e}_{j}$ denotes the $j$th unit vector in $\mathbb{R}^{r}$.
Case of $\bm{q}\neq\bm{0}$. Let $\bar{\bm{q}}=\bm{q}/\|\bm{q}\|_{1}$ for $\bm{q}$. From the equality $\bm{y}=W\bm{q}$, we obtain
$\|\bm{y}\|_{1}=\|W\bm{q}\|_{1}=\|\bm{q}\|_{1}\|W\bar{\bm{q}}\|_{1}\geq\rho(W)\|\bm{q}\|_{1}.$
Since $\rho(W)>0$, it follows that
$\|\bm{y}\|_{1}/\rho(W)\geq\|\bm{q}\|_{1}=\sum_{u=1}^{r}|\bm{q}(u)|\geq|\bm{p}(j)-1|.$
Therefore,
$\bm{p}(j)\geq 1-\|\bm{y}\|_{1}/\rho(W)$
for $j\in[r]$. Recall that $\mu(j)$ is defined as $\mu(j)=\max_{k\in\mathcal{K}}H(j,k)$ in (6). Since the columns of $H(\mathcal{K})$ are $\bm{h}_{k_{1}},\ldots,\bm{h}_{k_{s}}$, we can equivalently write
$\mu(j)=\max_{u=1,\ldots,s}\bm{h}_{k_{u}}(j).$
Using this representation, we observe that $\bm{p}(j)$ is upper-bounded by $\mu(j)$:
$\bm{p}(j)=\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{h}_{k_{u}}(j)\leq\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\mu(j)=\mu(j)$ (9)
where the last equality uses $\|\bar{\bm{c}}\|_{1}=1$ and $\bar{\bm{c}}\geq\bm{0}$. Hence,
$1-\mu(j)\leq\|\bm{y}\|_{1}/\rho(W)$ (10)
for $j\in[r]$. We now bound the $L_{1}$ norm of $\bm{y}$:
$\|\bm{y}\|_{1}$ $=\|\bm{v}_{i_{j}}-\sum_{u=1}^{s}\bar{\bm{c}}_{j}(u)\bm{v}_{k_{u}}-A(\mathcal{K})(\bm{c}_{j}-\bar{\bm{c}}_{j})\|_{1}$
$\leq\|\bm{v}_{i_{j}}\|_{1}+\sum_{u=1}^{s}|\bar{\bm{c}}_{j}(u)|\|\bm{v}_{k_{u}}\|_{1}+\|A(\mathcal{K})\|_{1}\|\bm{c}_{j}-\bar{\bm{c}}_{j}\|_{1}$
$\leq 2\epsilon+\|A(\mathcal{K})\|_{1}\|\bm{c}_{j}-\bar{\bm{c}}_{j}\|_{1}$
$\leq 2\epsilon+(1+\epsilon)\left|\|\bm{c}_{j}\|_{1}-1\right|.$ (11)
The last inequality uses
$\|A(\mathcal{K})\|_{1}=\|WH(\mathcal{K})+V(\mathcal{K})\|_{1}\leq\|W\|_{1}\|H(\mathcal{K})\|_{1}+\|V(\mathcal{K})\|_{1}\leq 1+\epsilon,$ (12)
from Assumption 1, and
$\|\bm{c}_{j}-\bar{\bm{c}}_{j}\|_{1}=\|\|\bm{c}_{j}\|_{1}\bar{\bm{c}}_{j}-\bar{\bm{c}}_{j}\|_{1}=\|\left(\|\bm{c}_{j}\|_{1}-1\right)\bar{\bm{c}}_{j}\|_{1}=\left|\|\bm{c}_{j}\|_{1}-1\right|\|\bar{\bm{c}}_{j}\|_{1}=\left|\|\bm{c}_{j}\|_{1}-1\right|.$
Combining inequalities (10) and (6.3) gives the desired result.
Case of $\bm{q}=\bm{0}$. In this case, we have $\bm{p}=\bm{e}_{j}$. Moreover, by (9), we have $\mu(j)\geq\bm{p}(j)$. Thus,
$1=\bm{p}(j)\leq\mu(j)\operatorname{\Leftrightarrow}1-\mu(j)\leq 0.$
Since
$\frac{2\epsilon+(1+\epsilon)\left|\|\bm{c}_{j}\|_{1}-1\right|}{\rho(W)}\geq 0,$
we obtain the desired result. ∎
Lemma 7.
Let the nearly $r$-separable matrix $A=M+V$ with $M=WH$ satisfy Assumption 1. Let $\mathcal{K}\in\Gamma(A)$. If $\epsilon<1$, then for each $j\in[r]$,
$\left|\|\bm{c}_{j}\|_{1}-1\right|\leq\frac{2\epsilon}{1-\epsilon},$
where $\bm{c}_{j}$ is defined in (8).
Proof.
From the representation $\bm{w}_{j}+\bm{v}_{i_{j}}=A(\mathcal{K})\bm{c}_{j}$ with $\bm{c}_{j}\geq\bm{0}$ in (8), we obtain the following bounds:
$1-\epsilon\leq\|\bm{w}_{j}+\bm{v}_{i_{j}}\|_{1}=\|A(\mathcal{K})\bm{c}_{j}\|_{1}\leq(1+\epsilon)\|\bm{c}_{j}\|_{1},$ (13)
$1+\epsilon\geq\|\bm{w}_{j}+\bm{v}_{i_{j}}\|_{1}=\|A(\mathcal{K})\bm{c}_{j}\|_{1}\geq(1-\epsilon)\|\bm{c}_{j}\|_{1}.$ (14)
We justify the bounds in (13). The left inequality follows from
$\|\bm{w}_{j}+\bm{v}_{i_{j}}\|_{1}\geq\|\bm{w}_{j}\|_{1}-\|\bm{v}_{i_{j}}\|_{1}\geq 1-\epsilon,$
and the right inequality follows from
$\|A(\mathcal{K})\bm{c}_{j}\|_{1}\leq\|A(\mathcal{K})\|_{1}\|\bm{c}_{j}\|_{1}\leq(1+\epsilon)\|\bm{c}_{j}\|_{1}$
where the final inequality uses (12) from the proof of Lemma 6. We justify the bounds in (14). The left inequality follows from
$\|\bm{w}_{j}+\bm{v}_{i_{j}}\|_{1}\leq\|\bm{w}_{j}\|_{1}+\|\bm{v}_{i_{j}}\|_{1}\leq 1+\epsilon,$
and the right inequality follows from
$\|A(\mathcal{K})\bm{c}_{j}\|_{1}$ $=\|(WH(\mathcal{K})+V(\mathcal{K}))\bm{c}_{j}\|_{1}$
$\geq\|WH(\mathcal{K})\bm{c}_{j}\|_{1}-\|V(\mathcal{K})\bm{c}_{j}\|_{1}$
$=\|\bm{c}_{j}\|_{1}-\|V(\mathcal{K})\bm{c}_{j}\|_{1}$
$\geq\|\bm{c}_{j}\|_{1}-\|V(\mathcal{K})\|_{1}\|\bm{c}_{j}\|_{1}$
$\geq(1-\epsilon)\|\bm{c}_{j}\|_{1}$
where the second equality follows from the nonnegativity of $W$, $H$, and $\bm{c}_{j}$, together with Assumption 1(a). Since $\epsilon<1$, inequalities (13) and (14) imply that
$\frac{1-\epsilon}{1+\epsilon}\leq\|\bm{c}_{j}\|_{1}\leq\frac{1+\epsilon}{1-\epsilon}\operatorname{\Leftrightarrow}\frac{-2\epsilon}{1+\epsilon}\leq\|\bm{c}_{j}\|_{1}-1\leq\frac{2\epsilon}{1-\epsilon}.$
Since
$0\leq\frac{2\epsilon}{1+\epsilon}\leq\frac{2\epsilon}{1-\epsilon},$
we conclude that
$\frac{-2\epsilon}{1-\epsilon}\leq\|\bm{c}_{j}\|_{1}-1\leq\frac{2\epsilon}{1-\epsilon}\operatorname{\Leftrightarrow}\left|\|\bm{c}_{j}\|_{1}-1\right|\leq\frac{2\epsilon}{1-\epsilon},$
completing the proof. ∎
We now prove Proposition 2.
(Proof of Proposition 2).
By definition, $\epsilon\geq 0$. Lemma 2, together with Assumption 1(a), implies that $\rho(W)\leq 1$. Since the proposition assumes that $\epsilon<\rho(W)/9$, we have
$0\leq\epsilon<\rho(W)/9\leq 1/9<1.$
Applying Lemmas 6 and 7 yields
$1-\mu(j)\leq\frac{4\epsilon}{\rho(W)(1-\epsilon)}$
for each $j\in[r]$. Moreover, since $\epsilon<\rho(W)/9\leq 1/9$, we obtain
$1-\mu(j)\leq\frac{4\epsilon}{\rho(W)(1-\epsilon)}<\frac{4\rho(W)/9}{8\rho(W)/9}=\frac{1}{2}.$
Hence, $1-\mu(j)<1/2$ for each $j\in[r]$. ∎
7 Concluding Remarks
We studied a data reduction technique that removes redundant pixels for endmember extraction from HSIs. This work was motivated by the high computational cost of self-dictionary methods when applied to large-scale HSIs. We conducted a theoretical analysis of this reduction step and developed a data reduction algorithm based on a splitting technique, called DRS. By combining DRS with a self-dictionary method that uses an LP formulation, we proposed a data-reduced self-dictionary method, REDIC, for endmember extraction. Numerical experiments demonstrated the effectiveness of REDIC in terms of both computational time and extraction accuracy.
We conclude this paper by suggesting directions for future research. As observed in Section 5.3, there are noticeable gaps between the MRSA distances of the outputs produced by DRS and the MRSA scores obtained by REDIC with $\lambda=0$, indicating that the extraction accuracy of REDIC still leaves room for improvement. One possible approach to addressing this limitation is to use a combinatorial method in place of the LP method within REDIC. Gillis [18] studied a combinatorial approach for a slightly generalized separable NMF problem and showed theoretically that it exhibits strong robustness to noise. This suggests that combinatorial methods may achieve superior endmember extraction accuracy. Due to their high computational cost, applying such methods directly to endmember extraction is impractical. However, by integrating them with DRS, it may be possible to achieve higher accuracy while maintaining reasonable computational time, potentially outperforming the current version of REDIC.
As mentioned in Section 2.1, the endmember extraction problem is closely related to problems such as topic modeling and community detection. It would be interesting to investigate whether the proposed data reduction technique can also be applied to these problems and thereby accelerate progress in these research areas.
Appendix: Analysis of the Outputs Produced by DR and DRS
We summarize the result concerning the outputs produced by DR and DRS in the following theorem.
Theorem 2.
Let $\mathcal{K}$ be the output produced by either DR or DRS for the input matrix $A$. Then the following statements hold:
  1. (a)
    $\mathcal{K}\in\Gamma(A)$.
  2. (b)
    For any $k\in\mathcal{K}$, the set $\mathcal{K}-k$ does not belong to $\Gamma(A)$.
To prove this theorem for the case of DRS, we establish the following lemma.
Lemma 8.
Let $\{\mathcal{I}_{1},\ldots,\mathcal{I}_{p}\}$ be a $p$-way partition of $[n]$. If $\mathcal{K}_{j}\subset\mathcal{I}_{j}$ satisfies $\operatorname{cone}(A(\mathcal{I}_{j}))=\operatorname{cone}(A(\mathcal{K}_{j}))$ for each $j\in[p]$, then $\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}))$.
Proof.
It suffices to show the inclusion $\operatorname{cone}(A)\subset\operatorname{cone}(A(\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}))$, since the reverse inclusion follows immediately from $n]\supset\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}$. Let $\bm{a}\in\operatorname{cone}(A)$. Then
$\bm{a}=\sum_{i\in[n]}\alpha_{i}\bm{a}_{i}$
for some coefficients $\alpha_{i}\geq 0$. Since $\{\mathcal{I}_{1},\ldots,\mathcal{I}_{p}\}$ is a $p$-way partition of $[n]$, this can be rewritten as
$\bm{a}=\sum_{j\in[p]}\sum_{i\in\mathcal{I}_{j}}\alpha_{i}\bm{a}_{i}.$
For each $j\in[p]$, define $\bm{b}_{j}=\sum_{i\in\mathcal{I}_{j}}\alpha_{i}\bm{a}_{i}$. Then $\bm{b}_{j}\in\operatorname{cone}(A(\mathcal{I}_{j}))=\operatorname{cone}(A(\mathcal{K}_{j}))$ by assumption. Thus, $\bm{b}_{j}$ can be written as
$\bm{b}_{j}=\sum_{k\in\mathcal{K}_{j}}\beta_{k}\bm{a}_{k}$
for some coefficients $\beta_{k}\geq 0$. Combining these, we obtain
$\bm{a}=\sum_{j\in[p]}\sum_{k\in\mathcal{K}_{j}}\beta_{k}\bm{a}_{k}.$
Because $\mathcal{I}_{j_{1}}\cap\mathcal{I}_{j_{2}}=\emptyset$ for $j_{1}\neq j_{2}$ and $\mathcal{K}_{j}\subset\mathcal{I}_{j}$, it follows that $\mathcal{K}_{j_{1}}\cap\mathcal{K}_{j_{2}}=\emptyset$ as well. Hence we can rewrite the above expression simply as
$\bm{a}=\sum_{k\in\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}}\beta_{k}\bm{a}_{k}.$
This proves the desired inclusion. ∎
DR iteratively constructs a finite number of sets $\mathcal{K}_{\ell}$. In the proof of Theorem 2, we denote by $t$ the total number of these sets. By construction, the sets $\mathcal{K}_{1},\ldots,\mathcal{K}_{t}$ satisfy
$[n]=\mathcal{K}_{1}\supset\mathcal{K}_{2}\supset\cdots\supset\mathcal{K}_{t}=\mathcal{K},$ (15)
where $n$ is the number of columns of $A$, and $\mathcal{K}$ is the output produced by DR. Furthermore, $\mathcal{K}_{\ell}$ and $\mathcal{K}_{\ell+1}$ satisfy
$\mathcal{K}_{\ell+1}=\mathcal{K}_{\ell}-i$ (16)
where the element $i\in\mathcal{K}_{\ell}$ is such that $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))$. We are now ready to prove Theorem 2.
(Proof of Theorem 2).
We first prove the statements for the case of DR, and then for the case of DRS.
Case of DR. We begin with statement (a). If $t=1$, then $n]=\mathcal{K}_{1}=\mathcal{K}_{t}=\mathcal{K}$, and hence $\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}))$, which establishes (a). Now assume $t\geq 2$. We show that $\operatorname{cone}(A(\mathcal{K}_{\ell}))=\operatorname{cone}(A(\mathcal{K}_{\ell+1}))$ for each $\ell=1,\ldots,t-1$, because this relation, together with $\mathcal{K}_{1}=[n]$ and $\mathcal{K}_{t}=\mathcal{K}$, implies statement (a). It suffices to prove the inclusion $\operatorname{cone}(A(\mathcal{K}_{\ell}))\subset\operatorname{cone}(A(\mathcal{K}_{\ell+1}))$, since the reverse inclusion follows from relation (15). Let $\bm{a}\in\operatorname{cone}(A(\mathcal{K}_{\ell}))$. Then
$\bm{a}=\sum_{k\in\mathcal{K}_{\ell}}\alpha_{k}\bm{a}_{k}$
for some coefficients $\alpha_{k}\geq 0$. By relation (16), we can rewrite this as
$\bm{a}=\sum_{k\in\mathcal{K}_{\ell+1}}\alpha_{k}\bm{a}_{k}+\alpha_{i}\bm{a}_{i}$
where $i$ is the element removed from $\mathcal{K}_{\ell}$ to form $\mathcal{K}_{\ell+1}$. Since $\bm{a}_{i}\in\operatorname{cone}(A(\mathcal{K}_{\ell}-i))=\operatorname{cone}(A(\mathcal{K}_{\ell+1}))$ by relation (16), we can express $\bm{a}_{i}=\sum_{k\in\mathcal{K}_{\ell+1}}\beta_{k}\bm{a}_{k}$ for some coefficients $\beta_{k}\geq 0$. Substituting this into the expression for $\bm{a}$ yields
$\bm{a}=\sum_{k\in\mathcal{K}_{\ell+1}}\alpha_{k}\bm{a}_{k}+\alpha_{i}\sum_{k\in\mathcal{K}_{\ell+1}}\beta_{k}\bm{a}_{k}=\sum_{k\in\mathcal{K}_{\ell+1}}(\alpha_{k}+\alpha_{i}\beta_{k})\bm{a}_{k}.$
Since each coefficient $\alpha_{k}+\alpha_{i}\beta_{k}$ is nonnegative, we conclude that $\bm{a}\in\operatorname{cone}(A(\mathcal{K}_{\ell+1}))$, proving the desired inclusion.
We next prove statement (b). Let $k\in\mathcal{K}$. Then, by the construction of the sequence $\mathcal{K}_{1},\mathcal{K}_{2},\ldots,\mathcal{K}_{t}$ generated by DR, there exists an index $\ell$ such that $k\in\mathcal{K}_{\ell}$ and $\bm{a}_{k}\notin\operatorname{cone}(A(\mathcal{K}_{\ell}-k))$. From relation (15), we have $\operatorname{cone}(A(\mathcal{K}-k))\subset\operatorname{cone}(A(\mathcal{K}_{\ell}-k))$. Thus, $\bm{a}_{k}\notin\operatorname{cone}(A(\mathcal{K}-k))$, even though $\bm{a}_{k}\in\operatorname{cone}(A)$. Hence, $\operatorname{cone}(A)\neq\operatorname{cone}(A(\mathcal{K}-k))$ for any $k\in\mathcal{K}$, which establishes statement (b).
Case of DRS. In this proof, we use the same notation as in the description of DRS (Algorithm 2), namely:
  • the $p$-way partition $\{\mathcal{I}_{1},\ldots,\mathcal{I}_{p}\}$ of $[n]$ produced in Step 1;
  • the output $\mathcal{J}_{u}$ produced by DR for the input matrix $A(\mathcal{I}_{u})$ in Step 2, for each $u\in[p]$;
  • $\mathcal{K}_{u}=\mathcal{I}_{u}(\mathcal{J}_{u})$ for $u\in[p]$;
  • $\mathcal{I}=\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}$;
  • the output $\mathcal{J}$ produced by DR for the input matrix $A(\mathcal{I})$ in Step 3.
Note that $\mathcal{K}$ is given by $\mathcal{K}=\mathcal{I}(\mathcal{J})$.
We first prove statement (a). By statement (a) for the case of DR, we have $\operatorname{cone}(A(\mathcal{I}_{u}))=\operatorname{cone}(A(\mathcal{K}_{u}))$ for each $u\in[p]$. Applying Lemma 8 then yields
$\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}_{1}\cup\cdots\cup\mathcal{K}_{p}))=\operatorname{cone}(A(\mathcal{I})).$
Furthermore, by statement (a) for DR, we obtain $\operatorname{cone}(A(\mathcal{I}))=\operatorname{cone}(A(\mathcal{K}))$. Consequently, $\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{K}))$, which proves statement (a) for DRS.
We next prove statement (b). From statement (b) for the case of DR, we have $\mathcal{J}-j\notin\Gamma(A(\mathcal{I}))$ for any $j\in\mathcal{J}$. This implies that
$\operatorname{cone}(A(\mathcal{I}(\mathcal{J})-i_{j}))\neq\operatorname{cone}(A(\mathcal{I})),$
where $i_{j}$ denotes the $j$th element of $\mathcal{I}=\{i_{1},i_{2},\ldots,i_{m}\}$ with $i_{1}<\cdots<i_{m}$. Note that $i_{j}$ belongs to $\mathcal{I}(\mathcal{J})$ because $j\in\mathcal{J}$. Recall that $\mathcal{K}=\mathcal{I}(\mathcal{J})$. Thus, by letting $k=i_{j}$, we obtain
$\operatorname{cone}(A(\mathcal{I}(\mathcal{J})-i_{j}))=\operatorname{cone}(A(\mathcal{K}-k)).$
Here, $k$ is an arbitrary element of $\mathcal{K}$, because $j$ is arbitrary in $\mathcal{J}$. Moreover, as shown in the proof of statement (a), we have $\operatorname{cone}(A)=\operatorname{cone}(A(\mathcal{I}))$. Consequently, $\operatorname{cone}(A(\mathcal{K}-k))\neq\operatorname{cone}(A)$ for any $k\in\mathcal{K}$. This proves statement (b) for the case of DRS. ∎